OpenStudy (anonymous):

Which of the following is the point-slope equation of the line with a slope = -5 and a point of (2, -1)? A. y - 1 = -5(x - 2) B. y + 1 = 5(x + 2) C. y + 1 = -5(x - 2) D. y - 1 = 5(x + 2)

OpenStudy (zale101):

Do you remember how the point slope equation looked like?

OpenStudy (anonymous):

i dont :(

OpenStudy (zale101):

\(\Large y-y_1=m(x-x_1)\) where m is the slope (x1, y1) is the given point

OpenStudy (zale101):

use what you were given and plug it into the equation

OpenStudy (haseeb96):

as (x1,y1)=(2,-1) amd m=-5 n now put them into the above equation of Zale101

OpenStudy (anonymous):

so y+1=-5x+15

OpenStudy (zale101):

how did you get 15?

OpenStudy (anonymous):

i plugged in 3 instead of 5 one sec

OpenStudy (anonymous):

so y+1=-5x-10

OpenStudy (zale101):

Oh okay, and by the way, you don't need to distribute because your choices aren't distributed. They just want you to plug it and get the signs right.

OpenStudy (anonymous):

so its choice c

OpenStudy (haseeb96):

yes it is C

OpenStudy (anonymous):

thank you both!

OpenStudy (haseeb96):

Good job @Zale101 and @princessS12

OpenStudy (zale101):

y+1=-5(x-2) y+1=-5x+10

OpenStudy (zale101):

No Problem!